Assume the following;

1. The interest rate in the Euro area is 4%;

2. The interest rate in Japan is 2%;

3. The current nominal exchange rate is 165 (€1 = 165 Yen);

4. Expected nominal exchange rate next year is 155.

Using the information above answer the following question:

a) What is the expected rate of appreciation or depreciation of the Euro? (precede the number with a + or minus to denote appreciation or depreciation, respectively, and use 2 decimal places - e.g +0.12 to mean an appreciation of 12%)?

b) How many Euro would a French resident expect to earn for each Euro invested in Japanese bonds for one year? (use 4 decimal places)?

c) How many Euro would a French resident expect to earn for each Euro invested in Japanese bonds for one year? (use 4 decimal places)?
Answer #1

Here , Current Exchange Rate = 165 Yen/ Euro

Interest Rate in Euro = 4%

Interest Rate in Japan = 2%

Expected Nominal Exchange Rate = 155Yen/ Euro

a) Forward Rate = Spot Rate * (1+ Foreign Rate)/ (1 + Domestic Rate)

= 165*(1+2%)/(1+4%) = 161.8269

Since, The forward rate is higher than expected rate, hence Euro is Expected to depreciate

Depreciation Rate = (161.8269 - 155) / 161.8269 = 4.22%

b)

1 Euro = 165 Yen

Now,

165 YEn after 1 year in Japan = 165*(1.02), since 2% is the interest rate in Japan

= 168.3 Yen

Now Expected Exchange Rate after 1 year 155 yen/ Euro

168.3 Yen = 168.3/ 155 Euro = 1.0858 Euro

Question c and b are same